OpenStudy (anonymous):

\[sin(x+iy)=sin(x)cos(iy)+sin(iy)cos(x)=sin(x)cosh(y)+isinh(y)cos(x)\] \[|sin(x+iy)|^2=?\] \[|sin(x+iy)|=\sqrt{sin^2xcosh^2y+sinh^2ycos^2x}\] \[|sin(x+iy)|^2=sin^2xcosh^2y+sinh^2ycos^2x\]

OpenStudy (anonymous):

How to get this to equal \[sin^2x+sinh^2y\]?

OpenStudy (anonymous):

\[\sin^2x(1-\cosh^2y)=\sinh^2y(\cos^2x-1)\]

OpenStudy (anonymous):

\[\cos^2x-1=-\sin^2x\] \[1-\cosh^2y=-\sinh^2y\] Yay!

hartnn (hartnn):

haha...u got it almost the same moment i got it..